OpenStudy (anonymous):

Write the following linear function in the form Ax+By=C. 1. y=3/4x-1 2. 2y=-5/3x+4

OpenStudy (anonymous):

\[y=\frac{ 3 }{ 4}x-1\]\[y+(\frac{ 3 }{ 4}x)=\frac{ 3 }{ 4}x=(-\frac{ 3 }{ 4}x)-1\]\[y-\frac{ 3 }{ 4}x=-1\]\[-\frac{ 3 }{ 4}x+y=-1\] \[\frac{ 3 }{ 4}x-y=1\] Can someone check my answer if correct.?????

OpenStudy (anonymous):

correct.. carry on...

OpenStudy (larseighner):

Yes, that is right. A, B, and C can be positive or negative, so as you have solved it A=3/4. B=-1, and C=1.

OpenStudy (anonymous):

\[2y=\frac{ -5 }{ 3}x + 4\] \[2y +\frac{ 5 }{ 3 }x=4\] \[5x + 6y = 12\]

OpenStudy (anonymous):

that would be for #2

OpenStudy (anonymous):

\[2y=-\frac{ 5}{ 3}x+4\]\[2y+(\frac{ 5 }{ 3}x)=-\frac{ 5 }{ 3}x+(\frac{ 5 }{ 3})+4\]\[2y-\frac{ 5 }{ 3}x=4\]\[-\frac{ 5 }{ 3}x+2y=4\] \[\frac{ 5 }{ 3}x-2y=-4\] help please. check my answer .?

OpenStudy (anonymous):

yes you are correct

OpenStudy (precal):

Yes but you need to multiply the whole thing by 3 to eliminate the fractions 5x-6y=-12

OpenStudy (anonymous):

can you do it for me . ? I don't get your point.???

OpenStudy (anonymous):

i already showed you a example

OpenStudy (anonymous):

I mean the whole solution . Can you do it ? please?

OpenStudy (anonymous):

no you are correct all the way were just telling you to multiply by 3 so you can cancel out the fraction.

OpenStudy (precal):

you distribute a 3 to each term 3( (5/3)x-2y)=3(-4) 5x-6y=-12 all of your work is correct but standard form is usually whole numbers for A and B

OpenStudy (anonymous):

as in multiply ? not in divide?

OpenStudy (precal):

distributive property is basic property in mathematics it means to multiply

OpenStudy (anonymous):

ahh. ok i will try.

OpenStudy (anonymous):

its pretty confusing in the beginning but it will get easier once you understand

OpenStudy (anonymous):

Ok i get it thanks...! :)

OpenStudy (anonymous):

So the final answer is 5x_6y+-12 ??

OpenStudy (anonymous):

5x-6y=-12 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no wait

OpenStudy (anonymous):

no its wrong

OpenStudy (anonymous):

huh. ?

OpenStudy (anonymous):

you made a mistake in your work

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

what is it?

OpenStudy (precal):

5x+6y=12

OpenStudy (anonymous):

well you did from this \[2y + (\frac{ 5 }{ 3 }x) = -\frac{ 5 }{ 3 }x+(\frac{ 5 }{ 3 }x) + 4\] to \[2y -\frac{ 5 }{ 3 }x = 4\]

OpenStudy (precal):

I think you changed a sign along the way

OpenStudy (anonymous):

and then so on

OpenStudy (anonymous):

but the answer is what precal said

OpenStudy (anonymous):

5x+6y=12 .is the final answer?

OpenStudy (anonymous):

thank you so much my friends , FriedRice and precal ... :)

OpenStudy (precal):

anytime